OpenStudy (anonymous):

Write a simplified expression for the area of the rectangle at the right. State all restrictions on a. (3a+9)/(2a-6) by (4a+4)/(a+3)

OpenStudy (campbell_st):

the values that a cannot take can be found by looking at the denominators in the dimensions you need to solve 2a - 6 = 0 and a + 3 = 0 these values would result in the denominators of each fraction being zero, which is undefined. so to simplify the area you need to multiply but first take out common factors \[\frac{3(a + 3)}{2(a - 3)} \times \frac{4(a + 1)}{(a + 3)}\] this will give \[\frac{3( a + 3)\times 4(a + 1)}{2(a - 3)\times(a + 3)}\] cancel any common factors... and then simplify hope this helps.

OpenStudy (anonymous):

i am still not quite sure how to simplify that.

OpenStudy (campbell_st):

well what can you see that is in the numerator and denominator..?

OpenStudy (anonymous):

a+3 cancel, but then what?

OpenStudy (campbell_st):

there is another common factor as well...its 2 so you have \[\frac{3\times2(a + 1)}{(a - 3)}\] can you simplify this... its just distributing the numerator

OpenStudy (anonymous):

but what happened to the 4 that was on the top

OpenStudy (anonymous):

of to get the 2 on the top it should have been 2(a+2) instead of 4(a+1)

OpenStudy (anonymous):

so its 3(a+2)/(a-3). is that simplified completely?

OpenStudy (anonymous):

What do you do next?? Gaahhh